First Announcement

Summer Workshop for Theoretical Physics

Hot Matter and Gauge Field Theories

2006, Rab, Croatia, 11-14 June 2006

Organized by Theoretical Physics Division of Rudjer Bošković Institute and Theoretical Physics Department of Oëtvös Lorand

University

Dear Colleagues,

It is our great pleasure to bring to your attention the first announcement of ”Hot Matter and Gauge Field Theories 2006”, the 1st International Workshop on Fundamental questions of Non-Abelian Quantum Gauge Theories and Cosmology, to be held in Rab, Croatia, during June 11-14, 2006. ”Rudjer Bošković” Institute, Zagreb, Croatia and Ëotvös Lorand University, Budapest, Hungary will cohost this conference.

Location:

The location of the conference will be the Dormitory of the ”Rudjer Bošković” Institute in Rab, Croatia. The conference includes four days of plenary talks.

The program of the conference will include the following topics:

Global and Collective Dynamics in A+A Collisions

Photons and Dileptons from Out of Equilibrium QCD

Renormalization in Coulomb gauge QCD

QCD at High Temperature and Density

Quark Matter in Astro-particle Physics

New Theoretical Developments

City and weather during the conference:

Rab (www.TZG-RAB.hr) is a small ancient town situated on the island of Rab, which is part of the Kvarner Bay archipelago in the north of the Croatian part of the Adriatic see. Some 9000 inhabitants are living there, mostly by agriculture, fishing, shipbuilding and tourism. The period of Workshop belongs to the early tourist season. The weather is typical mediterranian: dominantly sunny, warm to moderate hot, moderate sea temperatures.

How to reach Rab:

To reach Island of Rab from Zagreb, one has to choose highway A1 towards Karlovac, Senj, Jablanac; use ferry-boat Jablanac - Mišnjak (the journey lasts around 15min, and the frequency is 8-20 times a day) on the island of Rab, and finally by car to the town of Rab itself. In Croatia one is supposed to drive with 0% of alcohol in veins!

Hotels:

There is a posibility for accomodation in the Dormitory (where one can use the cooking facilities - outside of workshop time, or enjoy one of many available small restaurants around) or at some of Rab’s hotels (Carolina, Eva, Imperial, Padova, Barbat, Epario, Istra, …).

Contacts

E-mail: aandrasi@irb.hr (Andjelka Andraši) Tel: 385-1-4561134 Fax: 385-1-4680223 Web: http://xxx.2006 (to be available soon!)

Local Organizing Committee:

A. Andraši I. Andrić I. Dadić A. Patkos Z. Szep